40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 26 19:35:16 2019, Jon, Update, SUS, New c1susaux installed, passed first round of scripted testing IMG_3254.jpgIMG_3256.jpg
    Reply  Fri May 3 12:48:40 2019, gautam, Update, SUS, 1X4/1X5 cable admin D38CC485-1EB6-4B34-9EB1-2CB1E809A21A.jpeg
       Reply  Tue Jul 2 12:30:53 2019, gautam, Update, Electronics, Acromag crate switched to Sorensens 1X5Sorensens.pdfCDS_20190702.png
          Reply  Tue Jul 2 16:57:09 2019, gautam, Update, CDS, c1sus is flaky c1sus_timing.png
Message ID: 14718     Entry time: Tue Jul 2 12:30:53 2019     In reply to: 14592     Reply to this: 14719
Author: gautam 
Type: Update 
Category: Electronics 
Subject: Acromag crate switched to Sorensens 

[chub, gautam]

We crossed off another couple of bullets today.

It took me ~1 hour to realize that c1susaux requries the running of sudo /sbin/ifup eth0 to be run in order to see the martian network - why???

Activity:

  1. Stopped the c1susaux machine:
    • Moved alignment sliders of ITMX and ITMY to 0 as a precaution.
    • Shutdown the c1susaux machine so that it doesn't become unhappy with the missing Acromags when we power the unit down.
  2. Dialled down supply voltages on the +/- 15 V and +/- 20 V DC Sorensens. Current draw became 0 A on the front panel indicators.
  3. Chub tapped some new terminal blocks for +15 V DC and +20 V DC
    • This required some additional daisy chaining, which is why we dialled down the Sorensens.
    • New cables were made using the "standard" LIGO color scheme, which isn't really applicable in this case because we are using +15 V DC (orange sheath wire) and + 20 V DC (yellow sheath wire) whereas the closest LIGO standard voltages are +18 V DC and +24 V DC.
    • A test cable, presumably meant to be used in the electronics area (orange for +15 V DC) was destroyed for this work as we opted for speed rather than making a new cable.
  4. Disconnected bench power supplies that were powering the Acromags, and connected the new cables.
    • I opted to use 5 A fuses in the terminal blocks for these supplies as the current draw is pretty significant.
  5. Dialled the Sorensens back up to the nominal voltages:
    • Attachment #1 shows the front panels of the Sorensens before and after this work.
    • The current limit on the +20 V DC Sorensen had to be raised, because the Acromag box draws ~2.3 A on its own, whereas the previous current draw was 2.8 A.
  6. Brought the c1susaux machine back online. Took me a while to get to the bottom of why I wasn't able to see c1susaux on the martian, but eventually, I figured out the whole sbin/ifup thingy. 

I don't understand the exact chain of causation, but during this work, the fast c1sus model crashed. I had to go through a few iterations of the scripted vertex machine rebooting, but things seem to be back in a normal state now, see Attachment #2. Should probably run the IFO test suite to make sure everything is a-okay, but for now, I am able to lock the IMC so I'm moving on.

The main task remaining here is to take new pictures of everything and upload to the wiki. Also, need to update the Sorensen labels to reflect their current values, some of them are outdated.

Quote:
  • Take photos of the new setup, cabling.
  • Remove the old c1susaux crate from the rack to free up space, possibly put the PSL monitoring acromag chassis there.
  • Test that the OSEM PD whitening switching is working for all 8 vertex optics.(verified as of 5/3/19 5pm)
  • New 15V and 24V power cables with standard LIGO connectors need to be run from the Sorensenn supplies in 1X5. The chassis is currently powered by bench supplies sitting on a cart behind the rack.
  • All 24 new DB-37 signal cables need to be labeled.
  • New 96-pin DIN connectors need to be put on two ribbon cables (1Y5_80 B, 1Y5_81) in the 1X4 rack. We had to break these connectors to remove them from the back of the eurcrates.
  • General cleanup of any cables, etc. left around the rack. We cleaned up most things this evening.
  • Rename the host computer c1susaux2 --> c1susaux, and update the DNS lookup tables on chiara.
Attachment 1: 1X5Sorensens.pdf  2.394 MB  Uploaded Tue Jul 2 16:22:13 2019  | Hide | Hide all | Show all
1X5Sorensens.pdf
Attachment 2: CDS_20190702.png  21 kB  Uploaded Tue Jul 2 16:26:54 2019  | Show | Hide all | Show all
ELOG V3.1.3-